OpenStudy (anonymous):

A toddler's playground has two slides on either side. The slide on the left is situated at a 45° angle to the ground and is 52 centimeters away from the base of the playground. The slide on the right is positioned at a 30° angle to the ground and is 98 centimeters away from the playground's base. What is the total length of both slides on this playground rounded to the nearest tenth of a centimeter? Type the numeric answer only in the box below.

OpenStudy (anonymous):

Do you want me to add the picture it came with?

OpenStudy (campbell_st):

draw a picture  sin45 = 52/x sin 30 = 98/y x = 52/sin45 y = 98/sin30 hope this helps